# Calibrating Deep Neural Network using Euclidean Distance

### v1

@@ Abstract.
Probabilistic prediction, Focal Calibration Loss, Model Calibration
copyright: acmlicensedcopyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; isbn: ;submissionid: 123-A56-BU3-abcdccs: Computing methodologies Machine learning
Figure 1. Grad-CAM (Selvaraju et al., 2017) Heatmap Visualization of CheXNet Model (Rajpurkar et al., 2017) Predictions for Pneumonia Detection. This figure shows Grad-CAM heatmaps from the CheXNet model applied to two ChestX-ray14 test cases for pneumonia detection. The model predicts the probability of each thoracic disease and generates a corresponding pathology likelihood map. Subfigures (1.a, 2.a) show outputs from the model trained with cross-entropy loss; (1.b, 2.b) with focal loss; and (1.c, 2.c) with focal calibration loss, illustrating the impact of different loss functions on localizing disease-relevant regions.
[…]
TBC.
1. Introduction
@@
In web-based applications, ranging from search engines to medical imaging and diagnostic systems, well-calibrated models are essential for ensuring accurate and trustworthy outcomes. Deploying such models improves the reliability of predictions and fosters user confidence in automated decision-making, whether in image classifiers (Minderer et al., 2021) or large language models (Achiam et al., 2023). However, while Focal Loss, has proven effective in handling class imbalance, it often leads to overly confident predictions that are poorly calibrated (Lin et al., 2017). A well-calibrated predictor inherently aligns its predicted probabilities with the true likelihood of events, reducing the discrepancy between the model’s confidence and actual outcomes. This alignment minimizes the need for post-processing calibration methods, such as temperature scaling (Guo et al., 2017) or isotonic regression (Niculescu-Mizil and Caruana, 2005), which are typically employed to adjust miscalibrated outputs after model training. The post-processing gap is the difference between a model’s raw predictions and the calibrated probabilities required for reliable decision-making. Therefore, achieving a well-calibrated predictor directly within the loss function reduces this post-processing gap, leading to more accurate and trustworthy predictions without additional adjustments.
@@
Motivation and Contributions Recent studies have further established the theoretical foundation of focal loss in the context of classification calibration (Charoenphakdee et al., 2021). It has been demonstrated that focal loss is classification-calibrated, implying that a classifier trained with focal loss has the potential to reach the optimal Bayes classifier. Nonetheless, training with focal loss can yield classifiers that exhibit either underconfidence or overconfidence, making it an unreliable class-posterior probability estimator. Focal loss is not a proper scoring rule, which means it does not provide an unbiased estimation of class-posterior probabilities (Charoenphakdee et al., 2021). A proper scoring rule is designed to encourage the model to output probability estimates that match the true posterior distribution. Proper loss functions have been identified a local optimality condition necessary and sufficient for ensuring calibration (Blasiok et al., 2024). Certain proper loss functions, like Log-loss and quadratic loss, achieve calibration by ensuring that the loss cannot be significantly reduced through post-processing functions with specific Lipschitz continuity guarantees. This relationship highlights the optimization of proper loss functions, such as the Brier score (Brier, 1950), in achieving calibration in deep learning models, as suggested by Lakshminarayanan et al. (Lakshminarayanan et al., 2017). Given the prior established benefits of proper loss functions in calibration, this research investigates whether integrating a proper loss into the focal loss can boost its calibration as a class-posterior probability estimator. Our key contributions are as follows:
@@
  • We propose the Focal Calibration Loss (FCL), a theoretically grounded loss function that merges the traditional Focal Loss with a calibration loss term (proper loss), ensuring better probability calibration by uniquely aligning predicted probabilities with true class-posterior probabilities, thus effectively mitigating overconfidence and underconfidence.
@@
  • Through rigorous theoretical proofs, we demonstrate that minimizing the FCL yields classifiers with superior probability and classification calibration, resulting in a lower post-processing gap than with the Focal Loss alone.
@@
  • We empirically demonstrate that FCL significantly enhances model calibration and classification performance through extensive experiments on both in-distribution and out-of-distribution image and NLP datasets. Additionally, by employing the smooth calibration error (smCE) metric, we achieve more consistent calibration assessments, effectively addressing the discontinuity issues inherent in traditional bin-based metrics. FCL improves anomaly localization, recognition, and robustness in Chest X-ray imaging applications.
2. Preliminaries
2.1. Problem Formulation
@@
In multiclass classification, let $\mathcal{X}$ denote the input space and $\mathcal{Y}=\{1,2,\ldots,K\}$ the label space, where $K$ represents the number of classes. [^1] Vectors are denoted by boldface letters (e.g., $\bm{x}$), scalars by regular letters (e.g., $x$), and the $i$-th element of a vector $\bm{x}$ by $x_{i}$. The indicator function is $\mathbf{1}_{[\cdot]}$, and $\bm{x}^{\top}$ denotes the transpose of $\bm{x}$. Consider a dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{n}$ consisting of $n$ labeled instances independently drawn from an unknown probability distribution over $\mathcal{X}\times\mathcal{Y}$ with density $p(\bm{x},y)$. The goal is to find a classifier $f:\mathcal{X}\to\mathcal{Y}$ that minimizes the classification risk (Charoenphakdee et al., 2021):
(1) $\mathcal{R}^{\ell}(f)=\mathbb{E}_{(\bm{x},y)\sim p(\bm{x},y)}[\ell(f(\bm{x}),y)],$
@@
where $\ell(f(\bm{x}),y)=\mathbf{1}_{[f(\bm{x})\neq y]}$ is the zero-one loss, and $\ell\in[0,1]$. The true class-posterior probability vector is defined as $\bm{\eta}(\bm{x})=[\eta_{1}(\bm{x}),\ldots,\eta_{K}(\bm{x})]^{\top}$, where $\eta_{k}(\bm{x})=\mathbb{P}(y=k\mid\bm{x})$ is the true probability of class $k$ given input $\bm{x}$.
The Bayes-optimal classifier $f^{\ell,*}$, which minimizes the expected classification risk in Equation (1), is given by:
Definition 1 (Bayes-Optimal Classifier (Zhang, 2004)).
The Bayes-optimal solution for multiclass classification is defined as:
(2) $f^{\ell,*}(\bm{x})=\arg\max_{y\in\mathcal{Y}}\eta_{y}(\bm{x}).$
Knowing the true class-posterior probabilities $\bm{\eta}(\bm{x})$ allows for the determination of the Bayes-optimal classifier. However, the converse is not necessarily true (Bartlett et al., 2006).
2.2. Empirical Surrogate Risk
Training a neural network classifier typically involves learning a function $f:\mathcal{X}\rightarrow\Delta^{K}$ that maps an input $\bm{x}\in\mathcal{X}$ to a $K$-dimensional vector of predicted probabilities over the classes, denoted as $\hat{\bm{p}}(\bm{x})=f(\bm{x})=[\hat{p}_{1}(\bm{x}),\ldots,\hat{p}_{K}(\bm{x})]^{\top}$. Here, $\Delta^{K}$ denotes the $K$-dimensional probability simplex and define the predicted probability vector as:
(3) $\Delta^{K}=\left\{\bm{p}\in[0,1]^{K}\mid\sum_{k=1}^{K}p_{k}=1\right\}.$
where $\hat{p}_{k}(\bm{x})$ represents the predicted probability for class $k$. A common choice for $f$ is a deep neural network with a softmax output layer. For a given input $\bm{x}$, the decision rule $f^{*}$ selects the class with the highest predicted probability:
(4) $f^{*}(\bm{x})=\arg\max_{y\in\mathcal{Y}}\hat{p}_{y}(\bm{x}).$
Minimizing the classification risk in Equation (1) is impractical due to the finite sample size and the computational intractability of optimizing the zero-one loss because it is non-convex and non-differentiable (Bartlett et al., 2006; Zhang, 2004). Therefore, practitioners often minimize an empirical surrogate risk instead (Bartlett et al., 2006; Vapnik et al., 1998).
Definition 2 (Empirical Surrogate Risk (Vapnik et al., 1998)).
Let $\ell:\Delta^{K}\times\Delta^{K}\rightarrow\mathbb{R}$ be a surrogate loss function, and let $\mathbf{e}_{y_{i}}\in\{0,1\}^{K}$ be the one-hot encoded vector for class $y_{i}$. The empirical surrogate risk is defined as:
(5) ^$\hat{\mathcal{R}}^{\ell}(f)=\frac{1}{n}\sum_{i=1}^{n}\ell\big{(}f(\bm{x}_{i}),\mathbf{e}_{y_{i}}\big{)}.$
Applying regularization techniques helps mitigate overfitting (Ng, 2004). The choice of an appropriate surrogate loss function is crucial, as it affects the classifier’s performance. An ideal surrogate loss should be easier to optimize than the zero-one loss and ensure that minimizing the surrogate risk aligns with minimizing the expected classification risk (Charoenphakdee et al., 2021).
2.3. OC and UC Definitions
Definition 3 (Overconfidence and Underconfidence).
Let $f:\mathcal{X}\rightarrow\Delta^{K}$ be a classifier mapping inputs to predicted probabilities $\hat{\bm{p}}(\bm{x})=f(\bm{x})=[\hat{p}_{1}(\bm{x}),\ldots,\hat{p}_{K}(\bm{x})]^{\top}$, and $\bm{\eta}(\bm{x})=[\eta_{1}(\bm{x}),\ldots,\eta_{K}(\bm{x})]^{\top}$ denote the true class-posterior probabilities. The model $f$ is said to be:
@@ 2.3. OC and UC Definitions
Overconfident (OC)/Underconfident (UC) at $\bm{x}\in\mathcal{X}$ if the highest predicted probability exceeds/less than the highest true probability:
(6) $\delta_{\text{OC}}(\bm{x})$ $:=\max_{k\in\mathcal{Y}}\hat{p}_{k}(\bm{x})-\max_{k\in\mathcal{Y}}\eta_{k}(\bm{x})>0,$
(7) $\delta_{\text{UC}}(\bm{x})$ $:=\max_{k\in\mathcal{Y}}\hat{p}_{k}(\bm{x})-\max_{k\in\mathcal{Y}}\eta_{k}(\bm{x})<0.$
In other words, overconfidence occurs when the model assigns a higher probability to its most confident prediction than the true likelihood, potentially leading to overly assertive but incorrect decisions. Underconfidence occurs when the model is less certain than warranted, which may cause hesitation in decision-making.
2.4. Calibration Error
The Expected Calibration Error (ECE) is a widely used metric for assessing a model’s calibration. It quantifies the difference between predicted confidences and actual accuracies. Other metrics like Maximum Calibration Error (MCE), Adaptive-ECE (AdaECE), and Classwise-ECE provide additional insights. These metrics rely on binning strategies and are formally defined in Equations (18)–(22) (Appendix A.1). However, binning methods can suffer from issues like bin size selection and data sparsity within bins. To address these limitations, the Smooth Calibration Error (smCE) has been proposed, which avoids discrete binning by using continuous kernel density estimation (Błasiok and Nakkiran, 2023).
Definition 4 (Smooth Calibration Error (Blasiok et al., 2024)).
@@ 2.4. Calibration Error
Let $\mathcal{H}$ be the family of all 1-Lipschitz functions $\eta:[0,1]\to[-1,1]$. For predicted probabilities $\hat{\bm{p}}(\bm{x})=f(\bm{x})=(\hat{p}_{1}(\bm{x}),\hat{p}_{2}(\bm{x}),\ldots,\hat{p}_{K}(\bm{x}))$ and one-hot encoded true labels $\bm{y}=(y_{1},y_{2},\ldots,y_{K})$, smooth calibration error is:
(8) $\text{smCE}(f):=\sup_{\eta\in\mathcal{H}}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}\left[\sum_{k=1}^{K}(y_{k}-\hat{p}_{k}(\bm{x}))\eta(\hat{p}_{k}(\bm{x}))\right].$
2.5. Calibration and Entropy
Figure 2. Impact of the focusing parameter $\gamma$ on Expected Calibration Error (ECE) for (a) ResNet-18, (b) MobileNetV2, and (c) EfficientNet-B0, all trained with focal loss on CIFAR-10. Increasing $\gamma$ leads to higher ECE, indicating poorer calibration.
[…]
Three models trained with increasing gamma
Focal loss can be expressed in terms of the Kullback-Leibler (KL) divergence with entropy adjustments (Mukhoti et al., 2020a), as the inequality:
@@ 2.5. Calibration and Entropy
(9) $\mathcal{L}_{\text{focal}}(f(\bm{x}),\bm{y})\geq\text{KL}\big{(}\bm{\eta}(\bm{x})\,\big{\|}\,f(\bm{x})\big{)}+\mathbb{H}[\bm{\eta}(\bm{x})]-\gamma\mathbb{H}[f(\bm{x})],$
where $\text{KL}(\bm{\eta}(\bm{x})\,\|\,f(\bm{x}))$ is the KL divergence from the true distribution $\bm{\eta}(\bm{x})$ to the predicted distribution $f(\bm{x})$, and $\mathbb{H}[\cdot]$ denotes the entropy function (see Appendix A.9).
Calibration ensures that predicted probabilities reflect true outcome likelihoods. A well-calibrated model’s predicted probability $\hat{p}_{k}(\bm{x})$ matches the actual frequency of the outcome for class $k$. Focal loss adjusts the importance of examples using the focusing parameter $\gamma$. In Equation (9), the term $\text{KL}(\bm{\eta}(\bm{x})\,\|\,f(\bm{x}))$ encourages $f(\bm{x})$ to approximate $\bm{\eta}(\bm{x})$, promoting calibration. The term $\mathbb{H}[\bm{\eta}(\bm{x})]$ is constant and does not affect optimization.
For low $\gamma$, focal loss favors high-entropy (less confident) predictions, aiding calibration by avoiding overconfidence. However, as $\gamma$ increases, focal loss penalizes high-entropy predictions more, potentially leading to overconfident predictions that misrepresent true likelihoods, resulting in poor calibration, as illustrated in Figure 2.
2.6. Post-Processing Gap
We define the post-processing gap as the difference between the expected calibration loss $\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}[\mathcal{L}_{\text{calib}}(f(\bm{x}),\bm{y})]$ and the minimal expected calibration loss achievable through post-processing $f(\bm{x})\mapsto\kappa(f(\bm{x}))$ for some 1-Lipschitz function $\kappa$.
Definition 5 (Post-Processing Gap (Blasiok et al., 2024)).
Let $\mathcal{K}$ be the family of all post-processing functions $\kappa:[0,1]\rightarrow[0,1]$ such that the function $\delta(v)=\kappa(v)-v$ is 1-Lipschitz continuous. For a predictor $f:\mathcal{X}\rightarrow[0,1]^{K}$ and a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$, the post-processing gap $\text{pGap}_{\mathcal{D}}(f)$ is defined as:
$\text{pGap}_{\mathcal{D}}(f)=\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}\left[\mathcal{L}(f(\bm{x}),y)\right]-\inf_{\kappa\in\mathcal{K}}\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}\left[\mathcal{L}(\kappa(f(\bm{x})),y)\right].$
This gap quantifies the maximum improvement in calibration achievable through optimal post-processing. It serves as a benchmark to assess how close a model’s calibration is to the best possible calibration achievable via smooth transformations.
3. Proposed Method
3.1. Focal Calibration Loss

### v2

# Towards Trustworthy Deep Neural Network through Euclidean Distance

@@ Abstract
Machine Learning, ICML
1 Introduction
@@
In critical applications such as medical imaging and diagnostic systems, well-calibrated models are paramount for ensuring accurate and trustworthy outcomes. By aligning predicted probabilities with the true likelihood of events, well-calibrated predictors reduce the discrepancy between a model’s confidence and actual results, fostering user trust and minimizing the need for post-processing calibration methods (e.g., temperature scaling (Guo et al., 2017) or isotonic regression (Niculescu-Mizil & Caruana, 2005)). However, while Focal Loss (Lin et al., 2017) effectively addresses class imbalance, it often leads to overly confident predictions (Charoenphakdee et al., 2021). This shortfall highlights the post-processing gap—the difference between raw model outputs and the properly calibrated probabilities required for reliable decision-making.
Figure 1: Grad-CAM (Selvaraju et al., 2017) Heatmap Visualization of CheXNet Model (Rajpurkar et al., 2017) Predictions for Pneumonia Detection. This figure shows Grad-CAM heatmaps from the CheXNet model applied to two ChestX-ray14 test cases for pneumonia detection. The model predicts the probability of each thoracic disease and generates a corresponding pathology likelihood map. Subfigures (1.a, 2.a) show outputs from the model trained with cross-entropy loss; (1.b, 2.b) with focal loss; and (1.c, 2.c) with focal calibration loss, illustrating the impact of different loss functions on localizing disease-relevant regions.
@@
Motivation and Contributions. Recent studies have further established focal loss’s theoretical foundation in classification calibration (Charoenphakdee et al., 2021), showing that a classifier trained with focal loss can achieve the optimal Bayes classifier in theory. Nevertheless, focal loss is not a proper scoring rule, and training with it can yield underconfident or overconfident estimators—unreliable for class-posterior probability estimation. Proper scoring rules, such as the Brier score (Brier, 1950) or log-loss, encourage models to produce unbiased class-posterior probabilities (Lakshminarayanan et al., 2017) and ensure that post-processing does not significantly reduce the loss (Blasiok et al., 2024). Building on these insights, our work investigates whether integrating a proper loss into focal loss can improve its calibration as a class-posterior probability estimator. Our key contributions are:
@@
  • Focal Calibration Loss (FCL): We propose a new loss function that combines the traditional Focal Loss with a proper calibration term, uniquely aligning predicted probabilities with true class-posterior probabilities to mitigate overconfidence and underconfidence.
@@
  • Theoretical Guarantees: We prove that minimizing FCL yields superior probability and classification calibration compared to Focal Loss alone, effectively reducing the post-processing gap.
@@
  • Empirical Validation: Through extensive experiments on both in-distribution and out-of-distribution image and NLP datasets, we demonstrate that FCL significantly enhances model calibration and classification performance by using the smooth calibration error further ensures calibration assessment. We also show that FCL improves anomaly localization, recognition, and robustness in Chest X-ray imaging applications.
2 Preliminaries
2.1 Problem Formulation
@@
In multiclass classification, let $\mathcal{X}$ be the input space and $\mathcal{Y}=\{1,2,\ldots,K\}$ the label space of $K$ classes. [^1] Vectors are in bold (e.g., $\boldsymbol{x}$), scalars are in regular font (e.g., $x$), and the $i$-th element of a vector $\boldsymbol{x}$ is $x_{i}$. The indicator function is $\mathbf{1}_{[\cdot]}$, and $\boldsymbol{x}^{\top}$ denotes the transpose of $\boldsymbol{x}$. We observe a dataset $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{n}$ drawn i.i.d. from an unknown distribution with density $p(\boldsymbol{x},y)$ on $\mathcal{X}\times\mathcal{Y}$. We seek a classifier $f:\mathcal{X}\to\mathcal{Y}$ that minimizes the classification risk (Charoenphakdee et al., 2021):
$\mathcal{R}^{\ell}(f)=\mathbb{E}_{(\boldsymbol{x},y)\sim p(\boldsymbol{x},y)}\bigl{[}\ell\bigl{(}f(\boldsymbol{x}),y\bigr{)}\bigr{]},$ (1)
@@
where $\ell(\cdot,\cdot)$ is the zero-one loss, $\ell(f(\boldsymbol{x}),y)=\mathbf{1}_{[f(\boldsymbol{x})\neq y]}$. Although the true class-posterior probability vector $\boldsymbol{\eta}(\boldsymbol{x})=\bigl{[}\eta_{1}(\boldsymbol{x}),\ldots,\eta_{K}(\boldsymbol{x})\bigr{]}^{\top},\text{where }\eta_{k}(\boldsymbol{x})=\mathbb{P}\bigl{(}y=k\,\mid\,\boldsymbol{x}\bigr{)},$ would allow direct identification of the Bayes-optimal classifier:
$f^{\ell,*}(\boldsymbol{x})=\arg\max_{y\in\mathcal{Y}}\eta_{y}(\boldsymbol{x}),$ (2)
in practice we do not observe $\boldsymbol{\eta}(\boldsymbol{x})$ but only empirical samples $(\boldsymbol{x}_{i},y_{i})$. Hence, $\boldsymbol{\eta}(\boldsymbol{x})$ is purely theoretical and not assumed to be known.
2.2 Empirical Surrogate Risk
Minimizing $\mathcal{R}^{\ell}(f)$ in Eq. (1) directly is intractable because the zero-one loss is non-convex and non-differentiable (Bartlett et al., 2006; Zhang, 2004). Instead, one typically adopts a differentiable surrogate loss. Let
$f(\boldsymbol{x})=\hat{\boldsymbol{p}}(\boldsymbol{x})=\bigl{[}\hat{p}_{1}(\boldsymbol{x}),\ldots,\hat{p}_{K}(\boldsymbol{x})\bigr{]}^{\top}\in\Delta^{K},$
where $\Delta^{K}:=\{\boldsymbol{p}\in[0,1]^{K}\mid\sum_{k=1}^{K}p_{k}=1\}$. A common practice is to train a neural network with a softmax layer to output these probabilities, then define its prediction as:
$f^{*}(\boldsymbol{x})=\arg\max_{y\in\mathcal{Y}}\hat{p}_{y}(\boldsymbol{x}).$ (3)
The empirical surrogate risk is  (Vapnik et al., 1998):
$\hat{\mathcal{R}}^{\ell}(f)=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl{(}f(\boldsymbol{x}_{i}),\mathbf{e}_{y_{i}}\bigr{)},$ (4)
where $\mathbf{e}_{y_{i}}\in\{0,1\}^{K}$ is the one-hot encoding of $y_{i}$. By choosing a suitable surrogate $\ell$, we obtain a differentiable training objective that, in theory, aligns with minimizing the zero-one loss (Bartlett et al., 2006; Vapnik et al., 1998).
2.3 Overconfidence (OC) and Underconfidence (UC)
Definition 2.1 (Overconfidence and Underconfidence).
@@ 2.3 Overconfidence (OC) and Underconfidence (UC)
Let $f(\boldsymbol{x})=\hat{\boldsymbol{p}}(\boldsymbol{x})$ be a model’s predicted probability vector, and let $\boldsymbol{\eta}(\boldsymbol{x})$ be the theoretical posterior. The model $f$ is said to be OC/UC at $\boldsymbol{x}\in\mathcal{X}$ if the highest predicted probability exceeds/less than the highest true probability:
$\delta_{\mathrm{OC}}(\boldsymbol{x})$ $:=\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\;>\;0,$ (5)
$\delta_{\mathrm{UC}}(\boldsymbol{x})$ $:=\max_{k}\hat{p}_{k}(\boldsymbol{x})-\max_{k}\eta_{k}(\boldsymbol{x})\;<\;0,$ (6)
2.4 Calibration Error
Calibration assesses whether the predicted probabilities align with empirical frequencies. Traditional binning-based metrics, such as Expected Calibration Error (ECE), can be sensitive to bin sizes. Therefore, an alternative is the Smooth Calibration Error (smCE) (Blasiok et al., 2024):
Definition 2.2 (Smooth Calibration Error (smCE) [^2] In real-world settings, the true posterior $\boldsymbol{\eta}(\boldsymbol{x})$ is unknown; hence calibration metrics must be estimated from empirical data $\{(\boldsymbol{x}_{i},y_{i})\}$ rather than from $\boldsymbol{\eta}(\boldsymbol{x})$.).
@@ 2.4 Calibration Error
Let $\mathcal{H}$ be the class of all 1-Lipschitz functions $\phi:[0,1]\to[-1,1]$. For $\hat{\boldsymbol{p}}(\boldsymbol{x})=(\hat{p}_{1}(\boldsymbol{x}),\ldots,\hat{p}_{K}(\boldsymbol{x}))$ and one-hot $\boldsymbol{y}=(y_{1},\ldots,y_{K})$, the smooth Calibration Error is defined as:
$\text{smCE}(f):=\sup_{\phi\in\mathcal{H}}\;\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}}\Bigl{[}\sum_{k=1}^{K}\bigl{(}y_{k}-\hat{p}_{k}(\boldsymbol{x})\bigr{)}\,\phi\bigl{(}\hat{p}_{k}(\boldsymbol{x})\bigr{)}\Bigr{]}.$ (7)
2.5 Focal Loss Bound
A popular surrogate is the focal loss (Lin et al., 2017), which addresses class imbalance by downweighting easy examples. In prior theoretical analyses (Mukhoti et al., 2020a; Charoenphakdee et al., 2021), focal loss can be related to the KL divergence from $\boldsymbol{\eta}(\boldsymbol{x})$ to $\hat{\boldsymbol{p}}(\boldsymbol{x})$, giving a bound of form:
@@ 2.5 Focal Loss Bound
$\mathcal{L}_{\text{focal}}\bigl{(}\hat{\boldsymbol{p}}(\boldsymbol{x}),\mathbf{e}_{y}\bigr{)}\geq\mathrm{KL}\bigl{(}\boldsymbol{\eta}(\boldsymbol{x})\,\big{\|}\,\hat{\boldsymbol{p}}(\boldsymbol{x})\bigr{)}+\mathbb{H}\bigl{[}\boldsymbol{\eta}(\boldsymbol{x})\bigr{]}-\gamma\,\mathbb{H}\bigl{[}\hat{\boldsymbol{p}}(\boldsymbol{x})\bigr{]},$ (8)
where $\mathbb{H}[\cdot]$ is the entropy function and $\gamma$ the focal “focusing” parameter. In this bound, $\boldsymbol{\eta}(\boldsymbol{x})$ appears only in a theoretical sense to illustrate how focal loss penalizes higher-entropy predictions more strongly as $\gamma$ increases. Importantly, no step in the focal-loss training algorithm relies on direct knowledge of $\boldsymbol{\eta}(\boldsymbol{x})$; the model is trained on empirical data using a focal-loss objective, while $\boldsymbol{\eta}(\boldsymbol{x})$ remains a theoretical concept.
2.6 Post-Processing Gap
Finally, the post-processing gap quantifies how much a smooth recalibration function $\kappa(\cdot)$ could improve a classifier’s calibration under a given loss. Formally:
Definition 2.3 (Post-Processing Gap [^3] A small post-processing gap suggests that the raw outputs of $f(\boldsymbol{x})$ are already near-optimal for calibration under smooth transformations. By contrast, a large gap indicates substantial room for improvement via recalibration techniques. (Blasiok et al., 2024)).
Let $\mathcal{K}$ be 1-Lipschitz mappings $\kappa:[0,1]\to[0,1]$. Define $\delta(v)=\kappa(v)-v$. For a chosen calibration loss $\mathcal{L}$ and predictor $f$, $\text{pGap}_{\mathcal{D}}(f)$ is defined as:
$\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}}\bigl{[}\mathcal{L}\bigl{(}f(\boldsymbol{x}),y\bigr{)}\bigr{]}\;-\;\inf_{\kappa\in\mathcal{K}}\mathbb{E}_{(\boldsymbol{x},y)\sim\mathcal{D}}\bigl{[}\mathcal{L}\bigl{(}\kappa(f(\boldsymbol{x})),y\bigr{)}\bigr{]}.$ (9)
Reconciling Theory and Practice.
Although $\boldsymbol{\eta}(\boldsymbol{x})$ appears in theoretical statements (e.g.Eqs. (8) and (1)), it is not assumed to be known or used in training. Its primary role is in analysis: establishing properties like Bayes optimality, bounding calibration error, or characterizing over-/underconfidence. The actual training and evaluation rely exclusively on $\{(\boldsymbol{x}_{i},y_{i})\}$. Consequently, there is no contradiction between the theoretical need for $\boldsymbol{\eta}(\boldsymbol{x})$ in proofs and the final empirical solution (e.g. the surrogate risk minimization) that depends only on observable data.
3 Proposed Method
3.1 Focal Calibration Loss